### v1

# CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners

Abstract
Knowledge Editing (KE) enables the modification of outdated or incorrect information in large language models (LLMs). While existing KE methods can update isolated facts, they struggle to generalize these updates to multi-hop reasoning tasks that depend on the modified knowledge. Through an analysis of reasoning circuits—the neural pathways LLMs use for knowledge-based inference, we observe that current layer-localized KE approaches, such as MEMIT and WISE, which edit only single or a few model layers, struggle to effectively incorporate updated information into these reasoning pathways. To address this limitation, we propose CaKE (Circuit-aware Knowledge Editing), a novel method that enables more effective integration of updated knowledge in LLMs. CaKE leverages strategically curated data, guided by our circuits-based analysis, that enforces the model to utilize the modified knowledge, stimulating the model to develop appropriate reasoning circuits for newly integrated knowledge. Experimental results show that CaKE enables more accurate and consistent use of updated knowledge across related reasoning tasks, leading to an average of 20% improvement in multi-hop reasoning accuracy on MQuAKE dataset compared to existing KE methods. We release the code and data in https://github.com/zjunlp/CaKE.
CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners
Yunzhi ${}^{\text{\Scorpio\Pisces}}$thanks: Work done during Yunzhi’s visit to UCLA., Jizhan ${}^{\text{\Scorpio}}$, Jia-Chen ${}^{\text{\Pisces}}$, Ningyu ${}^{\text{\Scorpio}}$22footnotemark: 2, Shumin ${}^{\text{\Cancer}}$, Huajun ${}^{\text{\Scorpio}}$thanks: Corresponding Authors, Nanyun ${}^{\text{\Pisces}}$22footnotemark: 2, ${}^{\text{\Scorpio}}$ Zhejiang University ${}^{\text{\Cancer}}$ National University of Singapore ${}^{\text{\Pisces}}$ University of California, Los Angeles {yyztodd,fangjizhan,zhangningyu,huajunsir}@zju.edu.cn gujc@ucla.edu,violetpeng@cs.ucla.edu,shumin@nus.edu.sg
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in diverse tasks Yang et al. (2024a); Dubey et al. (2024); OpenAI (2024); Guo et al. (2025), achieving performance that rivals or even exceeds human experts. However, their practical deployment faces some critical limitations: parametric knowledge remains static after pretraining, making it challenging to keep up with evolving real-world information; their propensity for hallucinations also undermines reliability. Knowledge editing (KE) has emerged as a promising solution to update the knowledge in models precisely Mitchell et al. (2021); Wang et al. (2024c); Jiang et al. (2025). Although existing KE methods achieve good results on simple factual updates Yao et al. (2023); Zhang et al. (2024b), they often exhibit fundamental limitations: edits propagate inconsistently through related knowledge structures and downstream reasoning tasks Cohen et al. (2024); Qin et al. (2024); Yao et al. (2023); excessive focus on surface-level pattern matching Hoelscher-Obermaier et al. (2023), and locality issues for other unrelated knowledge and general ability Gu et al. (2024); Gupta et al. (2024).
Figure 1: The current edit cannot propagate the new knowledge to the reasoning circuit for multi-hop reasoning. We propose a circuit-aware edit to improve the model’s multi-hop reasoning performance involving the updated knowledge.
Figure 2: An overview of our work.
Our work specifically addresses the poor performance of edited models in downstream reasoning tasks that involve the updated knowledge Zhong et al. (2023); Zhang et al. (2024d). Consider a representative case in Figure 1 : after editing ‘Eddie Mathews, citizenship, United States → United Kingdom’, models correctly answer direct queries but fail multi-hop reasoning like ‘The capital of the country that Eddie Mathews was a citizen of is?’ (still outputting ‘Washington D.C.’). Critically, this is not merely an editing artifact: vanilla LLMs often correctly answer single-hop questions while failing their multi-hop counterparts Yang et al. (2024c); Biran et al. (2024), suggesting deeper architectural limitations in knowledge utilization.
We trace these limitations to a misalignment between KE strategies and the inherent reasoning architectures of LLMs. To investigate this disconnect, we examine how LLMs leverage knowledge in downstream reasoning tasks. Recent analysis suggests that knowledge is not merely statically stored but dynamically activated through specialized circuits Yao et al. (2024); Biran et al. (2024); Yu et al. (2025). However, these analyses overlook the phenomenon of LLM failures in reasoning circuits and fail to explore the underlying causes. Our investigation (§2) delves deeper into reasoning circuits, analyzing their structure and identifying the reasons behind failures in multi-hop reasoning cases. Specifically, the multi-hop reasoning emerges from coordinated computing circuits: early layers handle the first hop, extracting the bridge entity at the end-token of the first hop. This bridge entity, along with second-hop relation information, is then routed to the last token position in the middle layers. Subsequently, later layers utilize this information at the last token position to complete the reasoning process (Figure 2 (a)). We then analyze the entity and relation information at the last token position in failed multi-hop reasoning cases. Our observations reveal that critical information either fails to be properly routed to the last token position, or exhibits a weak signal, preventing effective reasoning. This explains why current KE methods underperform (§3.1): they optimize for isolated parameter changes rather than circuit-level integration needed for compositional reasoning, as shown in Figure 2 (b).
To bridge this fundamental gap, we propose Circuit-aware Knowledge Editing (CaKE) in §3.2. By explicitly aligning edits with the LLM’s native reasoning architecture, CaKE transforms static knowledge updates into generalizable knowledge reasoning circuit—models that not only store static edited facts but also dynamically apply them in downstream reasoning tasks. Specifically, for the updated knowledge, we design circuit-aware tasks that require the LLM to leverage this new knowledge for latent reasoning, ensuring its integration across distinct segments of the reasoning circuit, as depicted in Figure 2(c). To prevent unintended data leakage, we construct these tasks using ad-hoc features Zhang et al. (2024c) that are temporarily associated with the entities, such as ‘Japan is colored green. The capital city of the country colored in green is’. We then guide LLMs to construct the reasoning circuits to utilize the updated knowledge by training with the curated data. Extensive experiments (§4) demonstrate the effectiveness of CaKE, outperforming existing knowledge editing methods on the multi-hop editing benchmarks MQuAKE on LLMs of different sizes, including LLAMA3-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Yang et al. (2024a), and LLAMA3-70B-Instruct.
2 Analyzing Reasoning Circuit in LLM
In this part, we integrate previous findings on multi-hop reasoning and circuit analysis Yao et al. (2024); Biran et al. (2024); Yang et al. (2024b); Wang et al. (2024e) and factual recall Merullo et al. (2024) to illustrate how language models use knowledge to tackle multi-hop reasoning tasks. Based on this, we can better understand the cause of failures for multi-hop reasoning.
2.1 Data Preparation
We employ the WikiData subset proposed by  Biran et al. (2024) and name it HoppingTooLate, which contains 82,021 two-hop queries. We denote each fact as a triplet $(e,r,e^{\prime})$, where $e$ is the head entity, $r$ is the relation, and $e^{\prime}$ is the tail entity. We view two-hop queries as $(e_{1},r_{1},e_{2})$ and $(e_{2},r_{2},e_{3})$, where $e_{1}$ is the source entity, $e_{2}$ is the bridge entity, and $e_{3}$ is the target entity. We focus on the latent reasoning framework to evaluate whether a model can output the expected answer $e_{3}$ directly given the composite query $(e_{1},r_{1},r_{2},?)$. For example, for the two facts (Eddie Mathews, country_citizenship, United States),(United States, capital, Washington D.C.), the composite query is (Eddie Mathews, country_citizenship, capital,?). We transform the question into the natural language expression: ‘The capital of the country that Eddie Mathews was a citizen of is?’. In addition, we follow HoppingTooLate and define $t_{1}$ as the last token of the first-hop prompt (e.g., ‘the country that Eddie Mathews was a citizen of) and $t_{2}$ as the last token of the whole two-hop prompt (e.g., ‘The capital of the country that Eddie Mathews was a citizen of is’).
2.2 Multi-hop Reasoning Circuit
Building on the insights from prior work Biran et al. (2024); Yao et al. (2024), we can define a structured circuit mechanism for multi-hop reasoning in transformer-based LLMs, as illustrated in Figure 2(a). The three distinct computational phases: 1) The model processes the initial relation $r_{1}$ and entity $e_{1}$, encoding the bridge entity $e_{2}$ in the final token position of the first prompt segment ($t_{1}$). 2) Critical features, including $e_{2}$ and the second relation, $r_{2}$ are transferred to the last token position $t_{2}$, preparing for final resolution. 3) The model computes the target $e_{3}$ by resolving $r_{2}$ and $e_{2}$, giving the result in the final token position.
Model Entity Patch Relation Patch
LLaMA3-8B-Ins. 85.35 56.20
Qwen2.5-7B-Ins. 97.29 55.40
Table 1: Activation Patching Success Rates (%).
Hence, based on the linearity theory Hernandez et al. (2024), multi-hop reasoning in LLM can be formalized as:
$F_{n}(F_{n-1}(e_{n-1},r_{n-1}),r_{n})$ (1)
Each function $F_{n-1}$ produces a bridge entity $e_{n}$ for subsequent computation, demonstrating how intermediate results propagate vertically through network layers.
Evaluation
To validate this circuit hypothesis, we conduct a causal analysis to determine whether modifying the variables in the function $F$ leads to corresponding changes in the output. Our intervention strategy focuses on the critical last token position for the second hop in Figure 9, where intermediate variables are stored. Specifically, we consider: 1).Entity Patching: Replacing the representation of $e_{2}$ with an alternative entity $e_{patch}$. For example, given the prompt ‘The official currency of the country where the Eiffel Tower is located is’, we substitute the representation of the bridge entity ‘France’ with ‘China’, expecting the output to change to ‘Renminbi’. 2).Relation Patching: Replacing the representation of $r_{2}$ with an alternative relation $r_{patch}$. For instance, given the same prompt, we substitute the representation of ‘official currency’ with ‘capital’, expecting the output to change to ‘Paris’. A successful patching (producing $F_{2}(e_{patch},r_{2})$ or $F_{2}(e_{2},r_{patch})$) would confirm the model’s reliance on these specific representations for reasoning. We conduct experiments on LLAMA3-8B-Instruct and Qwen2.5-7B-Instruct and employ PatchScope Ghandeharioun et al. (2024) for targeted activation patching (detailed in §B.1). Table 1 shows that in both the LLAMA3-8B and Qwen2.5-7B models, substituting variable representations at the last token position leads to corresponding behavioral changes, particularly in entity patching, where accuracy exceeds 80%. These results provide mechanistic evidence for the reasoning circuit we identified before.
Model Metric Correct Inconsistent Incorrect
Cases Layer Cases Layer Cases Layer
LLAMA3 $e_{2}$ from $t_{1}$ 63.1% 6.3 75.2% 6.0 48.7% 8.2
$e_{2}$ from $t_{2}$ 67.8% 13.2 59.8% 9.8 17.7% 21.1
$r_{2}$ from $t_{2}$ 66.9% 14.0 49.0% 13.8 28.1% 13.7
$e_{3}$ from $t_{2}$ 56.5% 18.8 22.7% 20.7 18.3% 18.0
Qwen2.5 $e_{2}$ from $t_{1}$ 71.2% 4.3 74.1% 4.7 46.7% 5.1
$e_{2}$ from $t_{2}$ 52.9% 7.9 63.7% 9.5 18.9% 13.5
$r_{2}$ from $t_{2}$ 75.8% 8.1 75.2 % 10.4 44.8% 9.7
$e_{3}$ from $t_{2}$ 71.2% 16.4 39.4% 17.4 25.2% 11.4
Table 2: The results of LLAMA3-8B-Instruct (32 layers) and Qwen2.5-7B-Instruct (28 layers). Cases are the percentage of data we can detect the information, and Layer is the mean of the first layer in which we detect the information.
2.3 Circuit in Failure Phenomena
Then, we aim to understand why language models sometimes fail at multi-hop reasoning despite successfully answering individual single-hop questions. For instance, a model may correctly answer ‘the capital of Russia’ with ‘Moscow’ and ‘the country of citizenship of Fyodor Dostoyevsky’ with ‘Russia’, yet fail to answer the multi-hop question ‘the capital of the country of citizenship of Fyodor Dostoyevsky is’ correctly. To systematically analyze this issue, we focus on the second hop of reasoning, as the model typically performs well on the first hop. We categorize the data from the HoppingTooLate dataset [^1] We filter out short-cut cases as done by Biran et al. (2024). into three subsets based on the model’s behavior: Correct: The model answers both single-hop questions $(e_{1},r_{1},e_{2})$ and $(e_{2},r_{2},e_{3})$ correctly, as well as the multi-hop question $(e_{1},r_{1},r_{2},?)$. Inconsistent: The model answers both single-hop questions correctly but fails on the multi-hop question. However, some cases in the Correct set contain the bridge entity $e_{2}$ but have a different subject $e_{1}^{\prime}$ and relation $r_{1}^{\prime}$, for which the model correctly answers $(e_{1}^{\prime},r_{1}^{\prime},r_{2},?)$. This suggests that while the model can leverage knowledge in some contexts, it fails to generalize, indicating reasoning gaps rather than missing knowledge. Incorrect: The model answers both single-hop questions correctly but fails on the multi-hop question in all contexts $(e_{1}^{\prime},r_{1}^{\prime},e_{2})$. This implies a complete failure to employ the knowledge for multi-hop reasoning. To investigate these failure modes, we check whether the models construct the reasoning circuit by monitoring key variables ($e_{1}$, $e_{2}$, and $r_{2}$) at critical positions ($t_{1}$ and $t_{2}$) across the model’s layers using the PatchScope as Biran et al. (2024) did. Our analysis reveals several interesting patterns, extending beyond the ‘hopping too late’ problem identified by Biran et al. (2024).
We list the results in Table 2 (more details in Figure 8 in Appendix). For the correct subset, we observe strong evidence of the reasoning circuit functioning as expected: a large portion of $e_{2}$ is detected at both $t_{1}$ ($e_{2}$ from $t_{1}$) and $t_{2}$ ($e_{2}$ from $t_{2}$) in both LLAMA3 and Qwen2.5 models. The model correctly uses the $r_{2}$ and $e_{2}$ information at $t_{2}$ to produce the final answer $e_{3}$. Contrastly, in the Incosistent subsets, we can find that despite detecting $e_{2}$ and $r_{2}$ at $t_{2}$, the model often fails to produce the correct $e_{3}$ answer ($e_{3}$ from $t_{2}$: only 22.7% in LLAMA3 and 39.4% in Qwen2.5 of cases we can detect at $t_{2}$). We hypothesize that the $e_{2}$ information, though present, may be insufficient to trigger the second-hop reasoning circuit, leading to the failure to execute the function $F(e_{2},r_{2})$ effectively. What’s more, in the Incorrect subsets, we can find the needed $e_{2}$ information is rarely detected at the $t_{2}$ position ($e_{2}$ from $t_{2}$: Only 17.7% in LLAMA3 and 18.9% in Qwen2.5). Even when $e_{2}$ is detected, it typically emerges in much later layers (layer 21 in LLAMA3 and layer 13.5 in Qwen2.5), making it too late to be effectively utilized for the second-hop computation, aligned with Biran et al. (2024)’s findings. We conjecture the model fails to propagate $e_{2}$ to the $t_{2}$ position, resulting in the variable $e_{2}$ missing for conducting the $F(e_{2},r_{2})$ function.
Figure 3: Results of the intervention on the failure cases in multi-hop reasoning of LLAMA3 and Qwen2.5.
Evaluation
To test our hypothesis, we conduct interventions to enhance the information at the specific positions to see if we can improve the model’s performance in these failure cases. We test three ways: back-patching the $t_{1}$ and $t_{2}$ position as Biran et al. (2024) did, which would enhance the information at the position, and cross-position patching the information from $t_{1}$ to the $t_{2}$ position, which explicitly propagates the information from $t_{1}$ to $t_{2}$ (details in Figure 10 in Appendix). From the results in Figure 3, we can find a high success rate for all the inconsistent and incorrect cases, but they demonstrate different paradigms. For the inconsistent cases, back-patching would lead to better performance, while for the incorrect cases, patching knowledge from the $t_{1}$ to $t_{2}$ usually shows better outcomes. This proves our previous hypothesis that for the incorrect cases, due to the propagation failure, the model fails to move the $e_{2}$ to $t_{2}$ position, and manual routing via cross-patching can mitigate the issue. Meanwhile, for inconsistent cases, amplification via back-patching compensates the weak signal when valid $e_{2}$ representations reach $t_{2}$ but lack sufficient magnitude for subsequent reasoning.
3 Circuits-aware Knowledge Editing
3.1 Rethinking KE from the Circuit View
Despite the success of current KE on single facts benchmarks, from previous studies Zhang et al. (2024d); Zhong et al. (2023), and our analysis in §4, we can see that the edited model’s performance on multi-hop reasoning tasks is often unsatisfactory. Building on our previously identified circuit for multi-hop reasoning, we rethink the reason why current knowledge editing methods fail under multi-hop reasoning circumstances.
Unified Editing Details
When updating a piece of knowledge $(e,r,o\rightarrow o^{\prime})$, the most popular knowledge editing techniques would modify the parameters that are responsible for the knowledge. There are two kinds of paradigms: editing the Feed-Forward Networks (FFN) in the early layers, such as ROME Meng et al. (2022) and MEMIT Meng et al. (2023) or modifying the later layers’ FFN output, like WISE Wang et al. (2024c) and T-Patcher Huang et al. (2023). This is mainly based on the key-value memory features of the FFN Geva et al. (2020). However, some studies have queried the effectiveness of these localization settings Chang et al. (2024); Hase et al. (2024) as the localization area is not correlated to the performance of the knowledge editing methods. Here, we propose a unified view of the mechanisms and limitations from the circuit perspective. From Figure 2 (c), these two editing paradigms can be achieved by a gated function $\mathcal{G(\mathbf{x})}$:
$\text{FFN}_{\text{out}}(\mathbf{x})=\underbrace{\bm{W}\mathbf{x}}_{\text{Original term}}+\mathcal{G}(\mathbf{x})\cdot\underbrace{\bm{\delta}(\mathbf{x})}_{\text{Edit term}}$ (2)
$\mathcal{G}(\mathbf{x})=\begin{cases}1,&\mathbf{x}\in\bm{e}_{in}\\ 0,&\text{otherwise}\end{cases}$ (3)
When the gating function $\mathcal{G}(\mathbf{x})$ is activated by the input $\mathbf{x}$, the edit term $\bm{\delta}(\mathbf{x})$ is applied, thereby modifying the knowledge within the computational circuit. ROME-style would modify the weight $\bm{W}$ with a perturbation $\bm{\Delta}$ and obtain a new weight $\bm{W}^{\prime}=\bm{W}+\bm{\Delta}$. Here, $\bm{\delta}(\mathbf{x})=\bm{\Delta}\mathbf{x}$. When calculating the $\Delta$, ROME-style methods, apply the least squares estimation and null space constraint to make sure the $\Delta$ is only activated by the corresponding entity representation $\bm{e}_{in}$ and keep the original output for other representations. In parallel, WISE-style editing methods would directly introduce the new weight $\bm{W}^{\prime}$ that would be activated by the related representation $\bm{e}_{in}$, and $\bm{W}^{\prime}$ would encode the updated knowledge. Here, $\bm{\delta}(\mathbf{x})=(\bm{W}^{\prime}-\bm{W})\mathbf{x}$.
Figure 4: The target answer token’s rank in the vocabulary of different editing methods when editing the fact ‘The official language of Japan is Japanese $\rightarrow$ Korean.
Defect from circuit view
To see the editing mechanism better, we first compare the rank of the target answer at the last token position via MEMIT, WISE, and the original model in Figure 4. We edit the fact: ‘The official language of Japan is Japanese $\rightarrow$ Korean.’ From the figure, we can see that in the original model and MEMIT-edited model, the answer token is dealt with gradually through the mid-to-later layers, and MEMIT would make this happen in advance. The computation of the ROME-style method for the new knowledge recall is: $F(\tilde{e},r)=o^{\prime},\tilde{e}=e+\bm{\Delta}e$, which modifies the knowledge stored in the previous layer and gives us the new representation $e^{\prime}$ for further computing. On the contrary, the WISE method would directly alter the information at the edited layer as we can see the sharp drop at layer 29. In particular, the editing would take effect when the added or updated parameters are activated by query representation and work as $F(e,r)=o^{\prime}$.
Method Model MQUAKE-CF MQUAKE-CF-v2 MQUAKE-T
H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$
Pre-edited LLaMA3-8B-Ins 79.0 27.0 78.4 28.6 71.0 5.3
LoRA 66.0 27.6 64.7 24.6 92.3 66.0
WISE 38.2 24.0 37.2 21.0 63.5 62.9
MeLLo 16.5 16.1 19.5 16.0 42.3 50.1
ROME 86.8 17.6 86.4 15.5 89.5 8.4
MEMIT 76.3 11.5 74.0 10.0 86.0 3.7
AlphaEdit 66.1 10.1 63.7 8.5 73.4 1.0
IFMET 81.9 23.2 75.3 36.5 82.1 46.1
CaKE(ours) 90.6 57.3 90.1 57.1 91.5 81.4
Pre-edited 75.6 34.7 76.8 37.7 60.1 15.6
LoRA L-70B 93.1 53.2 90.5 50.2 90.1 90.6
MeLLo 8.0 6.4 8.6 9.9 11.6 32.9
CaKE(ours) 93.5 65.4 93.3 63.3 91.1 94.6
[…]
In single-hop knowledge editing, these kinds of methods would give us the correct information, but for the multi-hop cases, this would fail. As shown in Figure 2 (b), both these layer-specific editing methods cannot propagate the updated knowledge to the reasoning circuit, leading to unsatisfactory multi-hop reasoning performances. An essential requirement for these methods is that the gated function $\mathcal{G(\mathbf{x})}$ is activated by the specific representation $\bm{e}_{in}$. However, under the multi-hop reasoning scenario, the model would deal with different single-hop questions in different layers, like the two-hop reasoning circuit in §2: for ROME-style editing, if the new fact $(e,r,o\rightarrow o^{\prime})$ is the second-hop question and the entity $e$ appears after the edited layers, the gated function would $\mathcal{G(\mathbf{x})}$ not be activated and the model would still follow the previous stale knowledge $F(e,r)$ instead of $F(\tilde{e},r)$ and give us the wrong answer. Likewise, the WISE-style editing would retain reliance on the original knowledge when the new fact is finished in the former layers as the first hop, bypassing the edit function in later layers and cascading the error in the subsequent reasoning. Moreover, we conduct experiments to view the limitations better, and Table 8 show that editing the first-hop in WISE leads to poorer performance compared to editing the second-hop. Conversely, when using ROME to edit the second-hop, the performance is worse than when editing the first-hop. In conclusion, these layer-specific editing methods cannot learn the new knowledge generally to make the knowledge usable in downstream reasoning tasks.
3.2 Proposed Method: CaKE
Inspired by previous analysis, we propose a novel method, Circuit-aware Knowledge Editing (CaKE), which enhances the model’s ability to update and effectively utilize knowledge. CaKE comprises two key components: (1) generating circuit-aware training data that explicitly requires reasoning with the updated knowledge, and (2) training the model to construct robust reasoning circuits that integrate the new knowledge.
Data Generation
To ensure that the model builds effective reasoning circuits, we address two critical challenges: preventing failure propagation and mitigating weak signals (as identified in §2), while ensuring that updated knowledge is properly integrated across different layers (as described in §3.1).
Method Model MQUAKE-CF MQUAKE-CF-v2 MQUAKE-T
H-Acc.$\uparrow$ MAcc.$\uparrow$ H-Acc.$\uparrow$ MAcc.$\uparrow$ Hop-wise.$\uparrow$ MAcc.$\uparrow$
Pre-edited Qwen2.5-7B-Ins 73.4 40.7 72.8 39.5 56.1 15.6
LoRA 35.1 24.9 36.5 25.9 25.0 28.6
WISE 41.2 9.8 26.5 8.0 50.2 36.5
MeLLo 35.5 7.8 34.5 7.6 52.7 56.5
ROME 75.4 10.7 73.4 8.8 86.7 17.7
MEMIT 82.6 11.1 83.4 9.6 88.9 18.5
AlphaEdit 73.8 12.6 75.1 10.5 82.2 17.2
IFMET 83.7 25.7 84.6 24.5 90.0 52.8
CaKE(ours) 90.6 61.4 90.3 63.05 95.5 87.8
[…]
For each updated knowledge item, we construct the following contexts to mitigate these issues: (1) Original Narrative: We begin by generating straightforward factual statements that explicitly convey the updated information. For example, when updating the fact $k$: (PersonX, citizen_country, Switzerland → Japan), we use the narrative representation: ‘PersonX is a citizen of Japan’ and generate several paraphrases. These statements serve as the foundation for the model to learn the updated knowledge. (2) Circuit-aware Tasks: Next, we design specialized reasoning scenarios that address the identified circuit-level challenges, as illustrated in Figure 2(c). Moreover, to avoid introducing extraneous knowledge that could leak into downstream evaluations—and to test the generalization of our method (inspired by prior research Zhang et al. (2024c))—we incorporate ad-hoc features into these scenarios. These tasks link the facts with intermediate attributes or reasoning steps and fall into two categories: Late-layer Knowledge Integration: These tasks ensure that the updated knowledge is effectively learned in the later layers, alleviating issues such as weak signals and the limitations of ROME-style editing. For the fact $k$, we construct prompts like: ‘Suppose {random_entity_1} wears red clothes, {random_entity_2} wears blue clothes, and {PersonX} wears green clothes. The country of citizenship of the person in green is:’ Here, the model is expected to output ‘Japan,’ requiring it to employ the new fact $k$ in later layers. Reasoning Circuit Enhancement: These tasks require the model to use the updated knowledge for subsequent reasoning, thereby mitigating propagation failure, weak signal, and WISE-style’s limitations. Following the same fact $k$: ‘In a book about countries, Japan is mentioned on page 6 of the book, while China is mentioned on page 72. On which page of the book is the country of citizenship of the {PersonX} shown?’ Here, the model must first recall the updated citizenship (Japan) and then use this information to determine the correct page number (6). Furthermore, for each knowledge type, we develop specific task templates and leverage GLM-4-plus GLM et al. (2024) to generate data using randomly selected related entities (detailed in Appendix A).
Edit Training
After obtaining the curated circuit-aware data $\mathcal{D}$, we fine-tune the LLM using LoRA across all layers, enabling the model to optimize its internal knowledge organization. We minimize the cross-entropy loss $\mathcal{L}$ between the model’s outputs and the ground-truth tokens expressing the updated fact:
$\mathcal{L}=\mathbb{E}_{(\mathbf{x},\mathbf{y})\in\mathcal{D}}\left[-\sum_{t=1}^{|\mathbf{y}|}\log p(y_{t}\mid\mathbf{x},\theta_{\text{LoRA}})\right]$ (4)
where $\theta_{\text{LoRA}}$ represents the LoRA parameters, $\mathbf{x}$ is the input prompt, and $\mathbf{y}$ is the desired updated output sequence.
4 Experiments
4.1 Experiment Settings
We mainly utilize the multi-hop reasoning knowledge editing dataset MQuAKE Zhong et al. (2023), which considers different numbers of hops (from 2 to 4) and different positions of the knowledge used in the multi-hop questions. We utilize three versions of the datasets: MQuAKE-CF-3k and MQuAKE-CF-3k-v2, which are two subsets that contain different question types and editing hopping numbers, and MQuAKE-T is a time-aware knowledge editing benchmark.
Baselines and Models
We consider several knowledge editing baselines, including: IFMET Zhang et al. (2024d), AlphaEdit Fang et al. (2024), ROME Meng et al. (2022), MEMIT Meng et al. (2023),WISE Wang et al. (2024c) and MeLLo Zhong et al. (2023). Here, AlphaEdit, ROME, and MEMIT are methods that edit the model’s parameters at early layers; WISE adds additional parameters at later layers, and IFMET edits both the early and later layers’s FFN to achieve better multi-hop reasoning performance. MeLLo is a prompt-based retrieval-augmented method. We conduct experiments on LLAMA-3-8B-Instruct, Qwen-2.5-7B-Instruct, and LLAMA-3-70B-Instruct.
Evalutation Metric
Following Zhong et al. (2023), we evaluate model performance using Multi-hop Accuracy (MAcc) and Hop-wise Answering Accuracy (H-Acc). MAcc measures the accuracy of multi-hop question answering, while H-Acc assesses correctness at each reasoning step. Higher values indicate better performance. For KE, we also need to consider locality, which ensures edits do not affect unrelated knowledge and abilities. To assess this, we evaluate the model on general benchmarks, including CommonsenseQA Talmor et al. (2019), BigBenchHard Suzgun et al. (2023), MMLU Hendrycks et al. (2021), and GSM8k Cobbe et al. (2021).
4.2 Experiments Results
Main Results
Table 3 and Table 4 summarizes our results. Although current KE methods achieve high hop-wise accuracy (H-Acc.), their performance on the three versions of MQuAKE is quite low (with an average accuracy of less than 20%). For example, MEMIT and ROME achieve over 80% accuracy on single-hop questions in MQuAKE-v2; however, their accuracy on multi-hop reasoning drops to only around 10%, indicating that the LLM fails to effectively utilize the updated knowledge during reasoning. In contrast, CaKE demonstrates significant improvements in multi-hop reasoning. On the LLAMA3-8B-Instruct model, CaKE achieves accuracies of 57.3, 57.2, and 81.5 on MQuAKE-CF, MQuAKE-CF-v2, and MQuAKE-T, respectively—outperforming all compared methods. Additionally, IFMET, which also considers different layers for multi-hop reasoning but neglects the information flow within the circuit, performs not as well as CaKE. Moreover, when compared with RAG-based methods such as MeLLo, CaKE also yields better results. Furthermore, compared to the baseline LoRA tuning methods that simply incorporate the raw knowledge, the improvements observed with CaKE underscore the effectiveness of our approach. Interestingly, while LoRA demonstrates strong performance on the 70B model, reflecting enhanced learning capabilities in larger models, CaKE still achieves an additional improvement of approximately 10%. Results in Qwen-2.5-Instruct also demonstrate the same phenomenon.
Locality Performance
In this section, we evaluate the model’s performance on general ability benchmarks to ensure that acquiring new knowledge does not compromise its overall capabilities. As shown in Table 5, CaKE achieves performance comparable to the original model on both the LLAMA3-8B and Qwen2.5-7B models across different kinds of tasks, including math, commonsense, and diverse understanding tasks.
CSQA BBH MMLU GSM8k
LLaMA3-8B-Ins 76.09 67.89 63.83 75.20
MEMIT 76.08 67.88 63.82 75.21
ROME 72.98 61.37 62.95 74.59
CAKE 75.10 67.20 62.98 76.04
Qwen2.5-7B-Ins 82.31 33.39 71.80 82.26
MEMIT 82.39 37.37 71.80 81.96
ROME 72.57 34.22 63.38 72.21
CAKE 82.64 37.44 71.76 82.79
Table 5: Locality Performance on several general benchmarks of CaKE and other editing methods.
5 Analysis
5.1 Position and Number of Hop
We then examine the effects of the number of edits and the position of the updated knowledge in multi-hop scenarios, with results shown in Figure 5. Notably, even when the model is trained solely on two-hop questions, CaKE yields improvements across varying numbers of editing hops. The benefits are particularly pronounced for four-hop questions, where methods like IFMET (designed only for two-hop scenarios) struggle. Besides, CaKE enhances performance regardless of the position of the edited knowledge within the multi-hop questions, demonstrating the generalizability of CaKE .
5.2 Case Analysis
In this part, we show the cases in which the CaKE helps the model learn the multi-hop reasoning circuit and other methods fail. For illustration, we consider the two-hop question: ‘The capital city of the country that Eddie Mathews was a citizen of is’. Here, the editing case is (Eddie Mathews, citizenship, United States $\rightarrow$ United Kingdom), and the updated model is expected to output ‘London’. However, CaKE gives the correct answer, while other methods fail: MEMIT gives us the ‘Moscow’, AlphaEdit gives us ‘Birmingham’, and LoRA gives us ‘not known’. To further understand these differences, we analyze the computing circuit of each method to determine whether the updated model successfully propagates the bridge entity $e_{2}$ and relation $r_{2}$ to the last token $t_{2}$ position.
Figure 5: Accuracies of different number hops and edit-positions in MQuAKE-CF-3k-v2 on LLAMA3-8B-Instruct.
Figure 6 displays the logits of $e_{2}$ and $r_{2}$ at $t_{2}$ for models edited by different methods. As shown, the bridge entity $e_{2}$ in CaKE exhibits significantly stronger logits compared to those of AlphaEdit and MEMIT, indicating that CaKE effectively constructs the reasoning circuit and propagates the necessary information to the target position. Similarly, the $r_{2}$ information is more prominent in CaKE , further demonstrating its superiority in circuit construction and information flow.
5.3 Discussion with Chain-of-Thought
Figure 6: $e_{2}$ and $r_{2}$’s logits at $t_{2}$ in models after different knowledge editing methods.
Instead of directly providing an answer, chain-of-thought (CoT) reasoning generates intermediate steps sequentially. As proposed by Yang et al. (2024c), CoT not only facilitates knowledge activation in large language models but also transforms them into effective in-context reasoners. The CoT process builds a chain of relevant facts within the prompt context, where each step’s output serves as an in-context memory that subsequent steps can reference. This approach reduces the risk of losing track of intermediate facts as the sequence length increases, thereby promoting more coherent multi-hop reasoning. Moreover, because a significant portion of the model’s knowledge is stored in earlier layers, CoT can better leverage these neurons by decomposing complex questions into simpler sub-questions Wang et al. (2024g); Yao et al. (2025). Consequently, the reasoning circuit required for a single-hop inference is much simpler than that for multi-hop reasoning. This observation aligns with recent findings Li et al. (2024), which demonstrate that fast thinking without CoT leads to larger gradients and greater gradient disparities across layers compared to CoT. Nonetheless, inconsistencies in the intermediate reasoning steps still occur, highlighting potential areas for improvement. We believe that further analysis is needed to address these issues, and we leave this exploration for future work.
6 Related Work
Knowledge Learning and Editing
Knowledge editing Zhang et al. (2024b); Jiang et al. (2024a); Sun et al. (2024); Hsueh et al. (2024); Powell et al. (2024); Wang et al. (2024a); Rozner et al. (2024); Zhang et al. (2024a); Wang et al. (2024f); Shi et al. (2024); Huang et al. (2024); Guo et al. (2024); Wang et al. (2025b); Feng et al. (2025); Yang et al. (2025) has emerged as a promising approach for updating models in an ever-changing world. Current knowledge editing methods typically follow one of several strategies: modifying the MLP components in earlier layers Meng et al. (2022, 2023), enhancing the MLP in later layers Hartvigsen et al. (2023), or retrieving relevant facts as prompts Jiang et al. (2024b); Zhong et al. (2023).
Despite their potential, these methods often suffer from overfitting and performance collapse after sequential editing. To tackle these challenges, researchers Zhang et al. (2025); Wang et al. (2025a); Ma et al. (2024); Yang et al. (2024d) have identified that LLMs tend to focus excessively on the editing subject, and have proposed constraints to alleviate this issue. Additionally, several studies have sought to maintain model stability even after thousands of edits Cai and Cao (2024); Ma et al. (2025); Fang et al. (2024). However, most existing knowledge editing techniques concentrate on simple factual updates and frequently fail to generalize to more complex downstream tasks, such as multi-hop reasoning scenarios.
Model Interpretability
Knowledge editing is primarily based on the intrinsic knowledge mechanisms of neural models’ “black boxes” Ferrando et al. (2024). Consequently, understanding how knowledge in LLMs is acquired and stored has garnered significant attention Wang et al. (2024b). Recent studies Zhou et al. (2023) demonstrate that most knowledge is learned during the pretraining stage and is predominantly stored in the Feed-Forward Network Geva et al. (2020). Beyond these localized findings, researchers Geva et al. (2023); Yao et al. (2024) have investigated the computational circuits—the pathways connecting Transformer components—to elucidate how LLMs perform knowledge recall. Building on this, subsequent work has explored the relationship between knowledge editing and these circuits Ge et al. (2024). In contrast, our work focuses on the mechanisms underlying multi-hop reasoning in LLMs and aims to improve the generalization of edited knowledge.
7 Conclusion
We present CaKE, a framework designed to align knowledge editing with the inherent reasoning architectures of LLMs. By examining the multi-hop reasoning circuits within LLMs, we identify that existing knowledge editing methods fall short due to their isolated parameter adjustments, which fail to adequately propagate updated knowledge through the model’s reasoning circuit. CaKE addresses this gap by incorporating circuit-aware tasks that compel the model to dynamically integrate and utilize new knowledge during reasoning. Experimental results demonstrate that CaKE achieves generalizable multi-hop knowledge editing.
Limitation
Dataset
Our work primarily focuses on the factual knowledge embedded in large language models (LLMs) and their capacity for multi-hop reasoning over these facts. We recognize that LLM reasoning also encompasses other domains—such as long-form mathematics and reverse-curse reasoning—that merit further investigation.
Reasoning Pattern
As discussed in the previous analysis, we concentrate on direct reasoning phenomena. Current LLMs have shown impressive capabilities in slow-thinking paradigms, including chain-of-thought and reflective reasoning. Beyond direct reasoning, enhancing the utilization of knowledge within these paradigms represents an important avenue for future research.
Fine-grained Circuit Components
Our analysis revealed relational information within the circuits; however, CaKE currently does not delve deeply into these relationships. We believe that a more focused investigation into these components is necessary. Additionally, while our study emphasizes general circuit behavior, developing a more concise and effective method for knowledge editing remains an exciting challenge for future work.
Data Attribution
Although we demonstrate the ability to construct reasoning circuits using curated data, the connection between a model’s acquired abilities in its parameters and its training data is still underexplored. A deeper understanding of this relationship could lead to more efficient training processes and the generation of higher-quality synthetic data.
Appendix
Appendix A Setting Detail
Dataset
We list the details of the dataset in Table 6.
Model Correct Inconsistent Incorrect
LLaMA3-8B-Ins. 1,005 1,032 1,240
Qwen2.5-7B-Ins. 241 252 275
Table 6: The dataset we used in the analysis.
Environment Setting
We run our experiments on 2 NVIDIA-A800 GPUs. For data generation, we utilize glm-4-plus and glm-4-air and a total of 10,000,000 tokens (about 20 dollars) to generate all synthetic data for the whole dataset. We use LLM-Eval Gao et al. (2024) to test the model’s general performance.
Data Generation
We first construct the question template $\mathcal{T}$ for each relation type, and we list some of them in Table 7. We then generate the data using the following prompt:
Prompt for Constructing the circuit-aware data Here are some question templates for the specific relation. As you can see, the question use the knowledge in the input to conduct reasoning in different hops for multi-hop reasoning. Please generate 3 different questions based on the template. Please return a python json file. $\mathcal{T}$} Here is the input question:
Knowledge Type Template Answer
{target_person} works in the field of {target_field} . In a book related to different fields, Section A discusses {random_field}, Section B discusses {random_field}, and Section C discusses {target_field}. If you want to learn about {target_person}’s field, which section should you read? The working field of {target_person} is discussed in Section C.
In a biography book, Section A discusses the life of {random_person}, Section B discusses the life of {random_person}, and Section C discusses the life of {target_person}. The field of the person in Section C is? The person in Section C works in the field of {target_field}.
{target_person} speaks the language of {target_language}. The following facts are known: 1. {target_person} wears red clothes. 2. {random_person} wears blue clothes. 3. {random_person} wears green clothes. The language that the person in red clothes speaks is? The language that the person in red clothes speaks is {target_language}.
At a global company: {target_language}-speaking employees work in Team A. {random_language}-speaking employees work in Team B. In which team would {target_person} work when he/she is at work? {target_person} would work in Team A when he/she is at work.
Table 7: Sample templates for generating the circuit-aware data.
Appendix B Implementation Detail
B.1 Analyzing Method
Patch Scope
The process is carried out as follows. First, a source prompt, a source token, and a source layer are provided. The prompt is processed through the model’s forward computation, and the hidden representation $v$ of the source token at the specified layer is extracted and stored. This representation $v$ is the focus of our investigation, as we seek to determine whether it encodes a specific entity. Next, we employ the same prompt used by Ghandeharioun et al. (2024): “Syria: Syria is a country in the Middle East. Leonardo DiCaprio: Leonardo DiCaprio is an American actor. Samsung: Samsung is a South Korean multinational corporation. x” This prompt is passed through the model, but the hidden representation of ‘x’ is replaced with $v$ at a chosen target layer. The forward computation then proceeds, and the resulting generated text is analyzed to evaluate the effects of this substitution. We conduct different patch analyses and show them in Figure 9 and Figure 10. When we conduct back-patch and cross-patch, the source prompt and target prompt are the same.
Figure 7: The failure case of the multi-hop reasoning.
B.2 Editing Method
We utilize EasyEdit Wang et al. (2024d) to conduct our editing experiments. For ROME, MEMIT, WISE, AlphaEdit, and MeLLo, we directly employ the original parameters provided by their respective papers. Below, we introduce these methods in detail and describe our implementation.
ROME and MEMIT
ROME leverages causal analysis to identify knowledge within specific MLP layers and modifies the corresponding weight matrix using least squares approximation. It operates under the strong assumption that the MLP layers primarily store knowledge and injects new information into these layers iteratively using a Lagrangian remainder. In our experiments, we edit the 5th layer of both LLAMA3-8B-Instruct and Qwen2.5-7B-Instruct.
Similarly, MEMIT assumes that the FFN layers function as a knowledge key-value store. It directly modifies the parameters of selected layers through least squares approximation. Unlike ROME, which updates a single layer, MEMIT is a multi-layer editing algorithm capable of simultaneously updating hundreds or thousands of facts
IFMET
IFMET builds upon MEMIT by not only modifying earlier MLP layers in transformers but also adjusting later layers to enhance multi-hop reasoning for the edited knowledge. To ensure the updated knowledge propagates effectively, IFMET constructs an additional support set that reinforces learning in later layers. Based on our analysis in §2, we edit layers [17,18,19,20] for LLAMA3-8B-Instruct and layers [15,16,17,18] for Qwen2.5-7B-Instruct.
WISE
WISE represents a different approach to model editing, focusing on later layers instead of earlier ones. It modifies the model’s FFN output using a gating mechanism:
$\text{FFN}_{\text{out}}(\mathbf{x})=\begin{cases}\mathcal{G}(\mathbf{x})\cdot\mathbf{W}_{v^{\prime}}&\mbox{if $\mathcal{G}(\mathbf{x})>\epsilon$,}\\ \mathcal{G}(\mathbf{x})\cdot\mathbf{W}_{v}&\mbox{otherwise.}\end{cases}$ (5)
Here, $\mathcal{G}(\mathbf{x})$ is a gate function that computes the activation score of the hidden reprsentation: $\|\mathcal{A}(\mathbf{x})\cdot(\mathbf{W}_{v^{\prime}}-\mathbf{W}_{v})\|_{2}$. If the gate is activated, the model uses the updated knowledge to generate responses; otherwise, it relies on the original knowledge. Different methods define the gate function differently, but the core idea is to ensure that the updated memory aligns with relevant question representations.
Edit Method LLAMA3-8B Qwen2.5-7B
First_hop Second_hop First_hop Second_hop
ROME 16.66 7.81 10.57 8.33
WISE 49.85 67.36 8.33 33.59
Table 8: Performance comparison of edit methods across different positions for the edited fact.
MeLLo
MeLLo is a non-parametric editing method that modifies a model’s knowledge through prompting rather than weight updates. It maintains a memory of newly introduced facts and guides the model to decompose multi-hop queries into sub-questions. At each step, the model checks this memory to verify whether its existing knowledge contradicts the new facts. We follow the prompt structure provided in the original MeLLo method. However, in our experiments, we observe that the model struggles to consistently adhere to the intended reasoning pattern.
CaKE
We utilize the original LoRA Hu et al. (2022) and add parameters in the FFN module in the model. The hyperparameters are as follows:
  • epoch: [40, 50, 60]
  • batch size: [4]
  • learning rate: [1e-4]
  • rank: [8]
  • lora_alpha: [32]
Figure 8: The distribution of the layers allows us to detect the information from critical positions in the model via patch_scope.
Figure 9: The way we test the function of the second hop. If the model conducts the function at the later layers, changing the representation would change the output of the model.
Appendix C More Analysis
C.1 Concurrence or Reasoning?
Studies such as Yang et al. (2024c); Ju et al. (2024); Hou et al. (2023); Zhang et al. (2024c) those have discovered shortcuts in multi-hop reasoning. In the case of $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},e_{3}))$ (i.e., the query without r1), the model predicts correctly due to a high correlation between $e_{1}$ and $e_{3}$. For instance, given the query: “The capital city of the country where the Eiffel Tower is located is…” LLMs can sometimes provide the correct answer even without the intermediate context (‘the country where the Eiffel Tower is located’). In our analysis, we find that apart from the occurrence, the LLM would also sometimes conduct latent reasoning, such as ‘latently conducting the r1 completion’. If the model gives the correct e3 for $((e_{1},^{\prime\prime},e_{2}),(e_{2},r_{2},?))$ due to the occurrence, once we edit the $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$, the model would fail to give us the new answer. We select the shortcut data and conduct the editing in the first hop $(e_{1},r_{1},e_{2}\rightarrow e_{2}^{\prime})$ and then evaluate the model to see whether the edited model would output updated knowledge ($e_{1},r_{1},r_{2},e_{3}^{\prime}$). We conduct experiments on LLAMA3-8B-Instruct with the AlphaEdit method and demonstrate that about 65% percent of cases would give us the updated knowledge for the multi-hop questions, showing that edits to intermediate hops (e.g., updating the country) can disrupt reasoning when relying on pre-existing-shortcuts and correctly give us the newly updated reasoning results. This means that the LLM itself does not simply answer the questions due to the high correlation between e1 and e3, but actually conducts the latent reasoning.
C.2 Circuit Analysis
We present the model’s critical information detection results in Figure 8. The results indicate that knowledge is distributed across different layers, with incorrect cases appearing in later layers compared to correct and inconsistent cases. Determining the optimal layer for editing remains challenging, so we choose to adjust the model across all layers. In the future, we aim to refine our approach by performing more targeted edits.
C.3 Failure Phenomenon
In the multi-hop reasoning, we view several failure cases to see how the language model made mistakes for reasoning and we see it as the circuit competition. Here, we find the LLM tends to give us a wrong answer for the middle cases of the different entities that appeared in the middle steps. Take ‘The country that the creator of Hamlet was a citizen of’ as an example; the bridge entity here is ‘William Shakespeare’. We view ‘Hamlet’ as an entity that would influence the model to give us the results ‘Denmark’, which means the model has been distracted by other entities’ information. As shown in Figure 7, the model gives us the correct answer ‘England’ around layer 27 but output the wrong answer ‘Denmark’, which is actually the country of the ‘Hamlet’.
Figure 10: The way we conduct the backpatch and $e_{1}$ to $e_{2}$. We substitute the hidden representations from the source position to the target position.